OpenStudy (anonymous):

What is the side length of a cube with a volume of 0.008 m3? A. 0.000000512 m B. 0.002 m C. 0.02 m D. 0.2 m

OpenStudy (anonymous):

zerpdrix hurry

zepdrix (zepdrix):

Volume of a cube:\[\Large\bf\sf V\quad=\quad s^3\]Where s is the measure of each side. Our volume is given to be 0.008. So plug it in and solve for s.\[\Large\bf\sf 0.008\quad=\quad s^3\]Take the cube root of each side,\[\Large\bf\sf \sqrt[3]{0.008}\quad=\quad \sqrt[3]{s^3}\]The root and exponential function are inverse of one another, so on the right side it simplifies down,\[\Large\bf\sf \sqrt[3]{0.008}\quad=\quad s\]

zepdrix (zepdrix):

Are you allowed to use a calculator? That would be your next step if so.

OpenStudy (anonymous):

well no not really

OpenStudy (anonymous):

but whats the awnser